OpenStudy (anonymous):

square root of 36 over 5 6 over 5 6 over square root of 5 square root of 5 over 5 6 square root of 5 over 5

OpenStudy (phi):

\[ \frac{\sqrt{36}}{5} \] if you factor 36 into 6*6 this is the same as \[ \frac{\sqrt{6 \cdot 6}}{5} \] can you simplify ?

OpenStudy (anonymous):

sqrt(a/b)=sqrt(a)/sqrt(b) implies sqrt(36/5)=sqrt(36)/sqrt(5) therefore answer is 6/sqrt(5) sqrt_squareroot

OpenStudy (zzr0ck3r):

@giligails \[\frac{\sqrt{36}}{5}\]

OpenStudy (zzr0ck3r):

we know this because none of the answers have square roots in the denominator

OpenStudy (anonymous):

thanks guys